A recent study by the American Accounting Association revealed among 30 graduates in accounting 20 students graduating with a major in accounting select public accounting. Suppose we select a sample of 15 recent graduates.



A. What is the probability that exactly two select public accounting?



B. What is the probability that exactly five select public accounting?

1
Expert's answer
2022-05-06T12:42:56-0400

Let "X="the number of students who "X\sim Bin(n, p)"

Given "n=15, p=\dfrac{2}{3}, q=1-p=\dfrac{1}{3}"

A.


"P(X=2)=\dbinom{15}{2}(\dfrac{2}{3})^2(\dfrac{1}{3})^{15-2}"

"\approx0.000029"

B.


"P(X=5)=\dbinom{15}{5}(\dfrac{2}{3})^5(\dfrac{1}{3})^{15-5}"

"\approx0.006697"